# Dissecting Buchdahl’s limit: A surgeon’s guide to compact objects

### v1

@@ Abstract
One of the theoretical motivations behind the belief that black holes as described by general relativity exist in nature is that it is hard to find matter configurations that mimic their properties, especially their compactness. One of the classic results that goes in this direction is the so-called Buchdahl limit: a bound for the maximum compactness that, under a few assumptions, static fluid spheres in hydrostatic equilibrium can possibly have. We highlight two of the main assumptions that could be violated in physically realistic situations: i) isotropy and ii) an outward-decreasing monotonicity of the density profile, which, as we discuss in detail, can be understood as a form of energy condition. We construct a pair of toy models that exemplify how the Buchdahl limit can be overcome if any of these two assumptions is individually relaxed. In particular, we show that relaxing the monotonicity assumption alone yields a new, less restrictive compactness limit as long as the energy density is not allowed to become negative. If negative energies are permitted, the compactness of these toy models can be as close to the black hole limit as desired. We also discuss how these toy models represent some of the main features of realistic systems, and how they could be extended to find more refined models.
@@ I Introduction
Verifying whether the dark and compact objects that we observe in the universe strictly correspond to general relativistic black holes is still an open question that is attracting a large amount of interest and effort Collmar et al. (1998); Abramowicz et al. (2002); Carballo-Rubio et al. (2022). In an astrophysical context, any object that displays a sufficiently large gravitational mass in a sufficiently small region is believed to be a black hole. Probing and analyzing astrophysical black-hole systems could be our best chance to find a robust path beyond standard general relativity.
@@
Now, armed with the full Einstein equations one can analyze whether there could be, at least in principle, stellar objects more compact than neutron stars. Here comes the surprise. Under mild conditions there exists a limit to the maximum compactness attainable by a static, spherically symmetric sphere of perfect fluid matching the Schwarzschild metric at its surface. This so-called Buchdahl limit Buchdahl (1959); Bondi (1964); Wald (1984) is $C(R)<8/9$, and its origins can be traced back to Schwarzschild’s work on stellar interiors Schwarzschild (1916). Buchdahl proved that $C(R)<8/9$ as a theorem, making no strong hypothesis regarding the equation of state of matter beyond being barotropic. Its two central assumptions are just the isotropy of the fluid, namely that pressures in the tangential and radial directions are equal, and that the energy density is a positive, (outwards) monotonically decreasing function of the radius.
@@
Our aim in this paper is to explore the consequences and the physical viability of relaxing each individual assumption behind the Buchdahl theorem (assuming for simplicity a Schwarzschild exterior, i.e., vacuum, staticity and spherical symmetry). In particular, we are interested in understanding whether such relaxation leads to new, less restrictive, compactness bounds, or to no further compactness bounds at all. We restrict ourselves to static and spherically symmetric configurations, since potential generalizations to stationary and axisymmetric scenarios are far from straightforward.
@@
First, we will explore the role of the monotonicity and positivity of the energy density by introducing a stellar model composed of two thick layers of constant energy density: an internal core and an external crust, with the condition that the density is always larger for the crust than for the core. For this model, we analyze two very different situations: one in which we constrain the core density to be positive, and another in which we allow it to take negative values. We show that, in the first situation, a new bound on the compactness arises. We find numerically that this bound is approximately $C(R)\leq 0.9706$ (which is less restrictive than the Buchdahl bound $C(R)<8/9\simeq 0.888$, but still well-below the black hole compactness $C(R)=1$). Furthermore, we conjecture that a general bound on the compactness exists for a generic positive density independently of its behaviour, but at this stage we are unable to provide a complete proof. Meanwhile, when we allow for negative densities in the core, we show that there is no compactness bound; hence it is possible to build solutions as close to the black hole limit as desired, at least as long as we do not impose any energy condition on the fluid.
@@
The paper is structured as follows. In Section II, we revisit Buchdahl’s theorem as a warm-up exercise to show how the isotropy of the pressures and monotonicity of the positive density function enter into the proof. In Section III, we discuss the toy model that illustrates how to violate the Buchdahl limit relaxing the monotonicity of the density. In Subsec. III.1, we restrict ourselves to positive densities and show that there is an upper limit to the compactness that these objects can display, which is maximized when the internal core is “empty” (i.e., vanishing density in the internal layer). We conjecture that for positive monotonically increasing densities there should always exist an upper bound to the compactness. In Subsec. III.3, we consider the possibility of a negative inner core density, and we show that there is no limit to the compactness that these objects can have, illustrating the fact that allowing for any kind of negative density inner core would open the possibility of having objects as close to the black hole’s compactness as desired. In Section IV, we discuss the AdS star model: a particular case of bilayered star in which the inner core is described by AdS spacetime and the outer layer is reduced to a thin shell. In Section V, we introduce the model of a shell matching an inner Minkowskian core with an outer Schwarzschild metric. We show that it can get arbitrarily close to the compactness of a black hole, although there is a threshold beyond which energy conditions are violated. We use this result to illustrate the possibility of violating Buchdahl’s limit by considering anisotropic pressures. Finally, we conclude in Section VI with a discussion on the features of potentially realistic configurations that are reproduced by these toy models, and ways in which they could be improved in the future. Appendix A discusses the constant density perfect fluid spheres, and Appendix B provides the details of the construction of the AdS stars and the second toy model, i.e., the shell matching the internal Minkowskian or AdS core with the external Schwarzschild metric.
@@
II Buchdahl’s limit
We first review the Buchdahl limit in an attempt to present with clarity the steps in its derivation. The purpose of this section is twofold, serving both to settle the notation that will be used during this work, and to pose the problem that we will be studying. Importantly, we will enumerate the hypotheses behind the Buchdahl limit, i.e., the existence in general relativity of an upper bound to the compactness of stars in hydrostatic equilibrium.
@@ II Buchdahl’s limit
    The redshift function $f(r)$ is at least a $\mathcal{C}^{1}$ function while $h(r)$ at least $\mathcal{C}^{0}$. We integrate the equations from the centre $r=0$ with given pressures and densities $\{\rho_{c},p_{c}\}$ towards the surface $r=R$, defined by the condition $p(R)=0$. From that point on ($r>R$) we match an exterior (vacuum) Scwharzschild solution with $\rho=p=0$, independently of the value of $\rho\left(R^{-}\right)$.
@@
Proof: Demanding staticity automatically enforces $2M/R\leq 1$, since $C(R)=1$ will lead directly to a diverging pressure. This limit corresponds to a black hole, and physically it is not possible to have a matter content with an energy-momentum tensor that obeys energy conditions (in particular the dominant energy condition Wald (1984)) sitting statically onto an event horizon.
In this limit we have an event horizon, which, in virtue of Eq. (8), is incompatible with a gravitating perfect fluid as long as $\rho+p\neq 0$. Regular black hole models Bardeen (1968); Hayward (1996); Dymnikova (1992) satisfy $p+\rho=0$ and thus exhibit event horizons coexisting with perfect fluids. Leaving this particular case aside, it is possible to obtain a lower compactness limit by integrating Einstein equations with $\rho(r)\geq 0$ and $d\rho/dr\leq 0$ and requiring regularity of the solutions. Using the parametrization of the metric given by Eq. (1), Einstein equations read
$G_{tt}$ $=(rh^{2})^{-1}h^{\prime}+r^{-2}\left(1-h^{-1}\right)=8\pi\rho,$ (11)
$G_{rr}$ $=(rfh)^{-1}f^{\prime}-r^{-2}\left(1-h^{-1}\right)=8\pi p,$ (12)
$G_{\theta\theta}$ $=\frac{1}{2}(fh)^{-1/2}\frac{d}{dr}\left[(fh)^{-1/2}f^{\prime}\right]+\frac{1}{2}(rfh)^{-1}f^{\prime}-\frac{1}{2}(rh^{2})^{-1}h=8\pi p.$ (13)
@@
where we are using the respective $\mathcal{C}^{1}$ and $\mathcal{C}^{0}$ properties of $f$ and $h$ to match their values at the surface $r=R$ with their known Schwarzschild counterparts. We now multiply the equation by $rh^{1/2}$ and integrate inwards from the surface $r=R$ to the centre at $r=0$. We use the value of $f$ at the surface and the expression of $h$ in terms of $m(r)$ to find
@@
Observation 2: Notice that we are starting with a perfect fluid and hence the distribution of pressures is isotropic. As we illustrate below, the anisotropic case in which pressures in the angular directions are allowed to grow without bounds does not lead to any limit on the compactness of the object. By forcing the fluid to satisfy energy conditions, though, it becomes possible to find some compactness bounds as well Lindblom (1984); Ivanov (2002); Barraco et al. (2003); Boehmer and Harko (2006); Urbano and Veermäe (2019) (we will revisit this in Section V).
@@
Having detailed the hypotheses behind Buchdahl’s theorem, we devote the next three sections to presenting two stellar toy models for which said hypotheses do not hold, and to analyzing how this affects the maximum compactness they can possibly achieve.
@@ III Bilayered stars with non-monotonically decreasing density profiles
In this section we will focus on the study of a toy model displaying a very simple outward increasing density profile. Specifically, we consider a star whose internal structure displays two constant-density layers:
@@
Bilayered fluid stars satisfy Israel junction conditions at $R_{i}$ since, by virtue of (8), the discrete jump in $\rho$ between the inner and outer layers translates into a compensatory discontinuity in $p^{\prime}$ that guarantees the continuity of $\Phi^{\prime}$. Consequently, the metric components $g_{tt}$ and $g_{rr}$ are $C^{1}$ and $C^{0}$ functions, respectively. This discontinuous behaviour in the $tt$ component of the stress-energy tensor is of the same kind as the one that occurs at the surface of the star and, therefore, does not carry along the introduction of distributional stress-energy sources of any kind.
@@
We distinguish two situations: the case in which the inner density $\rho_{i}$ is strictly positive, which we explore in detail in Subsec. III.1, with the limiting case $\rho_{i}=0$ presented separately in Subsec. III.2; and the case in which the density of the inner layer is negative, which we discuss in Subsec. III.3. For positive inner densities, the Misner-Sharp function is everywhere positive, and we show that there is a maximum value of the compactness that can be reached around $2M/R\lesssim 0.9706$. For negative inner densities, regions with negative Misner-Sharp mass are allowed in the interior and we find that it allows for stars as compact as desired all the way up to the black hole limit.

### v2

# Beyond Buchdahl’s limit: bilayered stars and thin-shell configurations

@@ Abstract
One of the theoretical motivations behind the belief that black holes as described by general relativity exist in nature is that it is hard to find matter configurations that mimic their properties, especially their compactness. One of the classic results that goes in this direction is the so-called Buchdahl limit: a bound for the maximum compactness that spherically symmetric isotropic fluid spheres in hydrostatic equilibrium can possibly achieve with an outward-decreasing energy density. However, physically realistic situations could violate both isotropy and the monotonicity of the density profile. Notably, Bondi already showed that if the density profile is allowed to be arbitrary (but remains non-negative), a less restrictive compactness bound emerges. Furthermore, if negative energy densities are permitted, configurations can approach the black hole compactness limit arbitrarily closely. In this work we introduce a set of simple bilayered and thin-shell toy models designed to illustrate the effect of relaxing separately the assumptions of Buchdahl’s theorem. Within these models we highlight the existence of two special examples that we have called AdS stars and Einstein Static stars. We also discuss how these toy models may represent some of the main features of realistic systems, and how they could be extended to find more refined models.
@@ I Introduction
Verifying whether the dark and compact objects that we observe in the universe strictly correspond to general relativistic black holes is still an open question that is attracting a large amount of interest and effort Collmar et al. (1998); Abramowicz et al. (2002); Carballo-Rubio et al. (2022). In an astrophysical context, any object that displays a sufficiently large gravitational mass in a sufficiently small region is believed to be a black hole. Probing and analyzing astrophysical black hole systems could be our best chance to find a robust path beyond standard general relativity.
@@
Now, armed with the full Einstein equations one can analyze whether there could be, at least in principle, stellar objects more compact than neutron stars. Here comes the surprise. Under mild conditions there exists a limit to the maximum compactness attainable by a static, spherically symmetric sphere of perfect fluid matching the Schwarzschild metric at its surface. This so-called Buchdahl limit Buchdahl (1959); Bondi (1964); Wald (1984) is $C(R)<8/9$, and its origins can be traced back to Schwarzschild’s work on stellar interiors Schwarzschild (1916). Buchdahl proved that $C(R)<8/9$ as a theorem, making no strong hypothesis regarding the equation of state of matter beyond being barotropic. Its two central assumptions are just the isotropy of the fluid, namely that pressures in the tangential and radial directions are equal, and that the energy density is a positive, (outwards) monotonically decreasing function of the radius. The same bound seems to hold for the Einstein-Vlasov system Andréasson and Rein (2007).
Even when the density profile is not required to be a monotonically decreasing function of the radius, it is still possible to establish upper bounds on compactness, provided the density remains non-negative. Bondi Bondi (1964) found for this situation a general upper bound of $C(R)<12\sqrt{2}-16\approx 0.971$, which is less restrictive than the Buchdahl bound $C(R)<8/9\approx 0.889$, but still well below the black hole compactness $C(R)=1$ Bondi then proposed a model to illustrate how this limiting compactness could be attained. He vaguely describes this model as a core with no energy density joined to an external Schwarzschild geometry through a thin shell. However, he did not describe the characteristics of this thin shell. For instance, generically a thin shell obeying Israel’s junction conditions Israel (1966) contains tangential pressures which would introduce anisotropic pressures, against the very hypotheses of the set-up.
Bondi Bondi (1964) also demonstrated that if negative energy densities are permitted, the compactness can approach the black hole limit as closely as desired. Again, the specific model used to illustrate this result is based on a thin-shell construction that is not fully characterized. Less restrictive results but under more general conditions have also been reported in the literature. For instance, a bound of the form $12m(r)/r<$ 1 was found in Baumgarte and Rendall (1993) assuming positive energy density and isotropy (see Mars et al. (1996) for an interesting extension to the anisotropic case, where only regularity at $r=0$ and the condition $\rho+p_{r}+2p_{t}\geq 0$ are assumed). In Karageorgis and Stalker (2008), Karageorgis and Stalker present a variety of sharp bounds for the compactness of spherically symmetric configurations. On top of adding rigor to Bondi’s proof for the isotropic case, they also found other bounds by relaxing isotropy but imposing energy-like conditions.
@@
In this paper, we further explore the consequences and the physical viability of relaxing each individual assumption behind Buchdahl’s theorem within a spherically symmetric setup. In particular, we are interested in finding simple and realistic physical models that allow us to understand how relaxing these assumptions allows to bypass the limit. We restrict ourselves to static and spherically symmetric configurations, since potential generalizations to stationary and axisymmetric scenarios are far from straightforward.
@@
First, we will explore the role of the monotonicity and positivity of the energy density by introducing a stellar model composed of two thick layers of constant energy density: an internal core and an external crust, with the condition that the density is always larger for the crust than for the core. For this model, we analyze two very different situations: one in which we constrain the core density to be positive, and another in which we allow it to take negative values. We show that, in the first situation, we find that we are able to approach numerically Bondi’s bound $C(R)<12\sqrt{2}-16$ as much as desired. We will relate our models to Bondi’s, providing a clearer interpretation of his approach. We then relax the positivity condition by allowing for negative densities in the core. In this case, we show that it is possible to build solutions as close to the black hole limit as desired, at least as long as we do not impose any energy condition on the fluid, in agreement with Bondi’s results again.
@@
The paper is structured as follows. In Section II, we revisit key classical results from the literature. Subsection II.1 reviews Buchdahl’s theorem as a warm-up exercise, highlighting how the assumptions of isotropic pressure and monotonically-decreasing non-negative density function enter the proof. Readers familiar with the argument can skip this subsection. In Subsection II.2, we examine Bondi’s analysis, discussing the compactness limits he established for both non-negative and negative densities. In Section III, we discuss the toy model that illustrates how to violate the Buchdahl limit relaxing the monotonicity of the density. In Subsec. III.1, we restrict ourselves to positive densities, showing that the upper bound on compactness found by Bondi can be approached arbitrarily closely. We find that such limit is maximized when the internal core is “empty” (i.e., vanishing density in the internal layer) and study it in detail in Subsec. III.2. Furthermore, we compare our model with Bondi’s Model I. In Subsec. III.3, we consider the possibility of a negative inner core density, showing that any kind of negative density inner core would open the possibility of having objects as close to the black hole’s compactness as desired, in agreement with Bondi’s analyses. In Section IV, we discuss the AdS star model: a particular case of bilayered star in which the inner core is described by AdS spacetime and the outer layer is reduced to a thin shell. In Section V, we introduce the model of a shell matching an inner Minkowskian core with an outer Schwarzschild metric. We show that it can get arbitrarily close to the compactness of a black hole, although there is a threshold beyond which energy conditions are violated. We use this result to illustrate the possibility of violating Buchdahl’s limit by considering anisotropic pressures. Finally, we conclude in Section VI with a discussion on the features of potentially realistic configurations that are reproduced by these toy models, and ways in which they could be improved in the future. Appendix A discusses the constant density perfect fluid spheres, and Appendix B provides the details of the construction of the AdS stars and the second toy model, i.e., the shell matching the internal Minkowskian or AdS core with the external Schwarzschild metric.
@@
II Classic results: Buchdahl and Bondi
II.1 Buchdahl’s limit
We first review the Buchdahl limit in an attempt to present with clarity the steps in its derivation. The purpose of this section is twofold, serving both to settle the notation that will be used during this work, and to pose the problem that we will be studying. Importantly, we will enumerate the hypotheses behind the Buchdahl limit, i.e., the existence in general relativity of an upper bound to the compactness of stars in hydrostatic equilibrium. This subsection can be skipped by readers familiar with Buchdahl’s analysis.
@@ II.1 Buchdahl’s limit
    $f(r)$ is at least a piecewise $\mathcal{C}^{2}$ function, while $h(r)$ is at least piecewise $\mathcal{C}^{1}$. In particular, both $f^{\prime}$ and $h$ are continuous at $r=R$. We integrate the equations from the centre $r=0$ with given pressures and densities $\{\rho_{c},p_{c}\}$ towards the surface $r=R$, defined by the condition $p(R)=0$. From that point on ($r>R$) we match an exterior (vacuum) Scwharzschild solution with $\rho=p=0$, independently of the value of $\rho\left(R^{-}\right)$.
@@
Proof: Demanding staticity automatically enforces $2M/R\leq 1$, since $C(R)=1$ will lead directly to a diverging pressure. This limit corresponds to a black hole, and physically it is not possible to have a matter content with an energy-momentum tensor that obeys energy conditions (in particular the dominant energy condition Wald (1984)) sitting statically onto an event horizon. [^2] Regular black hole models Bardeen (1968); Hayward (1996); Dymnikova (1992) exhibit event horizons coexisting with anisotropic fluids, but these models fall outside the scope of this discussion. Leaving this particular case aside, it is possible to obtain a lower compactness limit by integrating Einstein equations with $\rho(r)\geq 0$ and $d\rho/dr\leq 0$ and requiring regularity of the solutions. Using the parametrization of the metric given by Eq. (1), Einstein equations read
$G^{t}_{\ t}$ $=(rh^{2})^{-1}h^{\prime}+r^{-2}\left(1-h^{-1}\right)=8\pi\rho,$ (11)
$G^{r}_{\ r}$ $=(rfh)^{-1}f^{\prime}-r^{-2}\left(1-h^{-1}\right)=8\pi p,$ (12)
$G^{\theta}_{\ \theta}$ $=\frac{1}{2}(fh)^{-1/2}\frac{d}{dr}\left[(fh)^{-1/2}f^{\prime}\right]+\frac{1}{2}(rfh)^{-1}f^{\prime}-\frac{1}{2}(rh^{2})^{-1}h=8\pi p.$ (13)
@@
where we are using the continuity of $f^{\prime}$ and $h$ to match their values at the surface $r=R$ with their known Schwarzschild counterparts. We now multiply the equation by $rh^{1/2}$ and integrate inwards from the surface $r=R$ to the centre at $r=0$. We use the value of $f$ at the surface and the expression of $h$ in terms of $m(r)$ to find
@@
Observation 2: Notice that we are starting with a perfect fluid and hence the distribution of pressures is isotropic. As we illustrate below, the anisotropic case in which pressures in the angular directions are allowed to grow without bounds does not lead to any limit on the compactness of the object. By forcing the fluid to satisfy energy conditions, though, it becomes possible to find some compactness bounds as well Lindblom (1984); Ivanov (2002); Barraco et al. (2003); Boehmer and Harko (2006); Andréasson (2008); Urbano and Veermäe (2019) (we will revisit this in Section V).
@@
II.2 Bondi’s Limit
In this section we review Bondi’s analysis of compactness bounds for isotropic stars Bondi (1964). We refer the reader to Karageorgis and Stalker (2008) for an ulterior analysis closing some technical details in Bondi’s proof. The framework of Bondi’s analysis is the one presented at the beginning of Section II.1. However, its key novelty lies in relaxing the assumption that the density profile must be monotonically decreasing.
Bondi’s analysis begins by introducing the variables $u(r)=m(r)/r$ and $v(r)=4\pi r^{2}p(r)$. In terms of these variables, the system reduces to:
$\frac{1}{r}\frac{dr}{du}=\frac{(1-2u)\frac{dv}{du}+u+v}{2v-(u^{2}+6uv+v^{2})},\quad 4\pi r^{2}\rho=u\frac{dv/du-\beta}{dv/du-\alpha},$ (22)
with
$\alpha=-\frac{u+v}{1-2u},\qquad\beta=-\frac{v}{u}\frac{2-5u-v}{1-2u}.$ (23)
Written in the $(u,v)$ variables, stellar solutions become trajectories in a $(u,v)$ diagram. From those trajectories, it is possible to obtain the physical variables immediately from Eq. (22), together with the equation for the redshift $\Phi$,
$\frac{1}{2}\,r\,\Phi^{\prime}=\frac{u+v}{1-2u}.$ (24)
Constant compactness curves, those described by $du/dr=0$, correspond in these variables to the hyperbolas
$H\equiv 2v-(u^{2}+6uv+v^{2})=0.$ (25)
Also, the integrals of $dv/du=\alpha$ are important and correspond to curves along which $dr/du=0$, i.e., there is a change in the compactness but the radius of the spheres is kept constant. They are described by the family of parabolas parametrized by $A$:
$P_{A}\equiv v=\sqrt{A(1-2u)}-1+u.$ (26)
Bondi identifies these parabolas with solutions describing thin shells since stellar solutions approaching these parabolas describe narrow matter layers with large densities.
Now, we need to consider how solutions describing a stellar interior behave in these $(u,v)$ variables. If we restrict to non-negative values of the density, $u$ needs to be non-negative as well. Furthermore, Bondi also shows that non-negative densities enforce non-negative pressures. Thus, an interior with non-negative densities is confined to lie within the first quadrant of the $(u,v)$ diagram in Fig. 1. In this figure we have drawn, in light blue, two curves representing stars with non-negative densities.
Figure 1: $(u,v)$ diagram of the kind used in Bondi’s proof Bondi (1964). Regular stellar solutions in the $u>0$ half-plane (with positive $m(r)$) describe curves that start from $(0,0)$ inside the $H>0$ region and increase until they cross $H=0$. After that, while moving within the $H<0$ region, they decrease, always with a more negative slope than $P_{A}$ curves, until they cross $v=0$. These solutions are always bound to take values below the dashed orange $P_{\rm A}$-curve with $A=9$, hence they must have a compactness $C(R)<C_{\rm B}$. Solutions in the $u<0$ half-plane (with negative $m(r)$) start from $(0,0)$ and can take negative $u$ values as large as desired. Then they have a turning point, cross $H=0$, and keep decreasing until they cross the $v=0$ axis. These solutions can have any compactness $C(R)<1$. We have drawn three representative curves meant to describe stars with increasing compactness.
In the $(u,v)$ diagram, a generic solution describing a stellar interior with finite central pressures (for example, the blue curves in Fig. 1) needs to be such that it starts from the origin $(0,0)$ and moves within region $H>0$, intersecting $P_{A}$ curves of increasing $A$ until it hits the hyperbola $H=0$ with exactly the same slope as the $P_{A}$ that is intersecting it at the crossing point (to ensure the finiteness of (22)).
Then, the solution enters the region $H<0$ and acquires negative slope, intersecting $P_{A}$ curves with decreasing $A$, eventually reaching $v=0$ at a given value of $u$ that fixes the compactness of the star. Now, we attempt to construct the most compact solution possible within the positive $(u,v)$ region. Restricting ourselves to $u\geq 0$, the highest point at which it is possible to cross $H=0$ is $v=2$. Once in the $H<0$ region, the curve that optimizes the compactness is the $A=9$ parabola, which intersects $v=0$ at $u=6\sqrt{2}-8$. Since a regular solution must lie below both these curves, this proves the existence of an upper bound for the compactness of stars with non-negative densities, namely
$C(R)<C_{\rm B}\equiv 12\sqrt{2}-16.$ (27)
The thick blue line in Fig. 1 represents a generic configuration near but below Bondi’s limit.
Allowing solutions to take values in the $u<0,v>0$ region, it is possible to obtain curves that grow in $v$ towards negative $u$ values and which are then matched with curves arbitrarily close to $P_{A}$ parabolas with $A$ values as large as desired. Since the $P_{A}$ curves cross $v=0$ at the positive $u$ values
$u_{+}=1-A+\sqrt{A\left(A-1\right)},$ (28)
we see that
$\lim_{A\to\infty}u_{+}=\frac{1}{2}-\frac{1}{8A}+\order{\frac{1}{A^{2}}},$ (29)
and hence we have that $C(R)<1$ and the black hole limit can be approached as much as desired. The darkest blue line in Fig. 1 represents a generic configuration with a negative energy core and a compactness beyond Buchdahl’s and Bondi’s bounds.
After presenting this general framework, Bondi proposed two specific stellar models, Model I and Model II. These models serve, respectively, to illustrate how to saturate the Bondi bound and how to construct configurations that surpass it thanks to negative energy cores. These models incorporate thin shells in their construction. At first glance, thin shells seem to require anisotropic pressures, which would be against the starting isotropy assumption. However, here we will show explicitly that these anisotropies do not arise since the pressures identically vanish due to a precise fine-tuning of the parameters of the models. We present a set of bilayered models with no distributional matter in the following sections as a complement to Bondi’s models. These models further illustrate how Buchdahl’s and Bondi’s limits can be approached and surpassed with specific examples.
@@ III Bilayered stars with non-monotonically decreasing density profiles
In this section, we will focus on the study of a toy model displaying a very simple outward increasing density profile. Specifically, we consider a star whose internal structure displays two constant-density layers:
@@
Bilayered fluid stars satisfy Israel junction conditions at $R_{i}$ since, by virtue of (8), the discrete jump in $\rho$ between the inner and outer layers translates into a compensatory discontinuity in $p^{\prime}$ that guarantees the continuity of $\Phi^{\prime}$. Consequently, the metric components $g_{tt}$ and $g_{rr}$ are piecewise $\mathcal{C}^{2}$ and $\mathcal{C}^{1}$ functions, respectively. This discontinuous behavior in the $tt$ component of the stress-energy tensor is of the same kind as the one that occurs at the surface of the star and, therefore, does not carry along the introduction of distributional stress-energy sources of any kind.
@@
We distinguish two situations: the case in which the inner density $\rho_{i}$ is strictly positive, which we explore in detail in Subsec. III.1, with the limiting case $\rho_{i}=0$ presented separately in Subsec. III.2; and the case in which the density of the inner layer is negative, which we discuss in Subsec. III.3. For positive inner densities, the Misner-Sharp mass function remains positive throughout the configuration, and we demonstrate that the maximum compactness bound found by Bondi, Eq. (27), can be approached arbitrarily. For negative inner densities, regions with negative Misner-Sharp mass appear in the interior, allowing for configurations that reach arbitrarily high compactness, all the way up to the black hole limit, in full agreement with Bondi’s conclusions. Unlike Bondi’s Model I and II, the bilayer model presented here gives rise to a well-defined geometry throughout.